OpenStudy (anonymous):

In a circle with radius 6, what is the measure (in degrees) of an arc whose length is 2π?

OpenStudy (dumbcow):

arc length = r*angle \[2 \pi = 6 \theta \] \[\theta = 2\pi/6 = \pi/3\]

OpenStudy (anonymous):

the arc length of a circle is the product of the angle it subtends and the radius meaning arc length = radius * angle thus angle = arc length/radius angle = 2*pi/6 angle = pi/3 or 60 degrees